OpenStudy (anonymous):

I need help with these 3 questions so I can ace my math test my brain is completely fried.... 1.Describe how the graph of y= x2 can be transformed to the graph of the given equation. y = x2 - 14 A. Shift the graph of y = x2 right 14 units. B. Shift the graph of y = x2 up 14 units. C. Shift the graph of y = x2 left 14 units. D. Shift the graph of y = x2 down 14 units.

OpenStudy (anonymous):

2. Describe how the graph of y= x2 can be transformed to the graph of the given equation. y = (x-14)2 - 9 A. Shift the graph of y = x2 down 14 units and then left 9 units. B. Shift the graph of y = x2 right 14 units and then up 9 units. C. Shift the graph of y = x2 right 14 units and then down 9 units. D.Shift the graph of y = x2 left 14 units and then down 9 units.

OpenStudy (solomonzelman):

erase this reply please

OpenStudy (solomonzelman):

\(f(x)+c\) is shifted c units up from \(f(x)\) \(f(x)-c\) is shifted c units down from \(f(x)\) (c is a positive number)

OpenStudy (anonymous):

3. Describe how to transform the graph of f into the graph of g. f(x) = x4 and g(x) = -x4 A Reflect the graph of f across the x-axis B Shift the graph of f down 1 unit. C Reflect the graph of f across the x-axis and then reflect across the y-axis. D Reflect the graph of f across the y-axis.

OpenStudy (solomonzelman):

compare \(y=x^2-14\) to \(y=x^2\). do you see the \(f(x)-c\) and \(f(x)\) thing here?

OpenStudy (solomonzelman):

okay, I am not doing more than 1 problem in the same post at a time. Have a nice day and hope you find someone who will help you.